# https://r4dl.github.io/CoMe/

### 25 Mar 2026 v1

@@ 3.1 Preliminaries
where $w_{i}(\bm{r})=\alpha_{i}(\bm{r})\prod_{j=1}^{i-1}(1-\alpha_{j}(\bm{r}))$. Following recent works [19, 55, 70], we evaluate the opacity $\alpha_{i}(\bm{r})$ directly in 3D at the point $\bm{x}^{*}_{i}$ of maximum contribution along the ray:
@@ 3.4 Variance Reducing Losses for Blending
While confidence-steered optimization already improves mesh quality, we still observe cases where view-dependent appearance is represented with geometry; most frequently, this is characterized by highly opaque primitives just behind surfaces, which are blocked in different amounts by semitransparent primitives depending on the viewing angle. This is in part enabled by the use of alpha-blending (see Eq.˜1): Since only the blended final color is used to compute the loss, individual primitives have no incentive to correctly model radiance.
Previous work [74] replaces the loss function of NeRF to apply the photometric loss to the individual color samples along the ray instead of the blended result, which effectively constrains such extreme cases. We find that naïvely applying this method to 3DGS significantly deteriorates reconstruction quality, because the vital D-SSIM loss term cannot be applied per-primitive. Instead, we introduce an additional weighted per-primitive $\mathcal{L}_{2}$-loss with the ground truth pixel color $\mathbf{I}$ to complement the standard 3DGS training loss:
@@
${\mathcal{L}}_{\text{normal-var}}=\sum_{i=0}^{N-1}w_{i}(\bm{r})\left\lVert\bm{n}_{i}-\bm{N}\right\lVert_{2}^{2}=1-\rVert\bm{N}\lVert_{2}^{2},$ (14)
@@ 0.A.2 Confidence
As $\hat{C}\rightarrow 0$, this quantity explodes; with our minimum confidence value, the gradient reaches at most 75-75 with $\beta=7.5\times 10^{-2}$. The maximum value was chosen to ensure that the gradients flowing back to individual Gaussians remain reasonably scaled, preventing further densification of already well-reconstructed regions.
@@ 0.C.1 Color Variance Loss
$\frac{\partial{\mathcal{L}}_{\text{color-var}}}{\partial\text{sh}(\bm{\theta}_{i},\bm{d})}=2w_{i}(\bm{r})(\text{sh}(\bm{\theta}_{i},\bm{d})-\mathbf{I})$ (22)
We can compute the gradient with regard to $\alpha_{i}$ as follows:
$\frac{\partial{\mathcal{L}}_{\text{color-var}}}{\partial w_{i}}$ $=\left\lVert\text{sh}(\bm{\theta}_{i},\bm{d})-\mathbf{I}\right\lVert_{2}^{2}$ (23)
$\frac{\partial w_{i}}{\partial\alpha_{i}}$ $=\prod_{j=1}^{i-1}(1-\alpha_{j}(\bm{r}))$ (24)
$\frac{\partial{\mathcal{L}}_{\text{color-var}}}{\partial\alpha_{i}}=\frac{\partial{\mathcal{L}}_{\text{color-var}}}{\partial w_{i}}\cdot\frac{\partial w_{i}}{\partial\alpha_{i}}$ $=\left\lVert\text{sh}(\bm{\theta}_{i},\bm{d})-\mathbf{I}\right\lVert_{2}^{2}\prod_{j=1}^{i-1}(1-\alpha_{j}(\bm{r}))$ (25)
This gradient does not account for the effect of $\alpha_{i}$ on the Gaussians $j>i$, which include $(1-\alpha_{i})$ in the weight term. Because $\alpha_{i}<1$ by definition, $\frac{\partial w_{j}}{\partial\alpha_{i}}=-\frac{w_{j}}{1-\alpha_{i}}$ holds and
$\sum_{j=i+1}^{N-1}\left\lVert\text{sh}(\bm{\theta}_{j},\bm{d})-\mathbf{I}\right\lVert_{2}^{2}\left(-\frac{w_{j}(\bm{r})}{1-\alpha_{i}(\bm{r})}\right)=-\frac{1}{1-\alpha_{i}}\sum_{j=i+1}^{N-1}w_{j}(\bm{r})\left\lVert\text{sh}(\bm{\theta}_{j},\bm{d})-\mathbf{I}\right\lVert_{2}^{2}$ (26)
@@
$\frac{\partial{\mathcal{L}}_{\text{color-var}}}{\partial\alpha_{i}}=\left\lVert\text{sh}(\bm{\theta}_{i},\bm{d})-\mathbf{I}\right\lVert_{2}^{2}\prod_{j=1}^{i-1}(1-\alpha_{j}(\bm{r}))-\frac{1}{1-\alpha_{i}}\sum_{j=i+1}^{N-1}w_{j}(\bm{r})\left\lVert\text{sh}(\bm{\theta}_{j},\bm{d})-\mathbf{I}\right\lVert_{2}^{2}$ (27)
For efficiency, $\sum_{j=i+1}^{N-1}w_{j}(\bm{r})\left\lVert\text{sh}(\bm{\theta}_{j},\bm{d})-\mathbf{I}\right\lVert_{2}^{2}$ is computed for each pixel during the forward pass. Then, we subtract the weighted squared difference for each primitive in turn to find the remaining variance for the next primitive. Note that as we build upon SOF [50], we inherit the front-to-back backward pass from StopThePop [49].
@@ 0.C.2 Normal Variance Loss
$=\underbrace{\sum_{i=0}^{N-1}w_{i}(\bm{r})}_{=1}-2\underbrace{\left(\sum_{i=0}^{N-1}w_{i}(\bm{r})\bm{n}_{i}\right)\cdot\bm{N}}_{=\left\lVert\bm{N}\right\lVert_{2}^{2}}+\left\lVert\bm{N}\right\lVert_{2}^{2}\underbrace{\sum_{i=0}^{N-1}w_{i}(\bm{r})}_{=1}$ (31)
$=1-\left\lVert\bm{N}\right\lVert_{2}^{2}$ (32)
@@
$\frac{\partial{\mathcal{L}}_{\text{normal-var}}}{\partial\alpha_{i}}=\left\lVert\bm{n}_{i}-\bm{N}\right\lVert_{2}^{2}\prod_{j=1}^{i-1}(1-\alpha_{j}(\bm{r}))-\frac{1}{1-\alpha_{i}}\sum_{j=i+1}^{N-1}w_{j}(\bm{r})\left\lVert\bm{n}_{i}-\bm{N}\right\lVert_{2}^{2}$ (34)

### 22 Apr 2026 v2

@@ 3.1 Preliminaries
where $w_{i}(\bm{r})=\alpha_{i}(\bm{r})\prod_{j=0}^{i-1}(1-\alpha_{j}(\bm{r}))$. Following recent works [20, 57, 72], we evaluate the opacity $\alpha_{i}(\bm{r})$ directly in 3D at the point $\bm{x}^{*}_{i}$ of maximum contribution along the ray:
@@ 3.4 Variance Reducing Losses for Blending
While confidence-steered optimization already improves mesh quality, we still observe cases where view-dependent appearance is represented with geometry; most frequently, this is characterized by highly opaque primitives just behind surfaces, which are blocked in different amounts by semi-transparent primitives depending on the viewing angle. This is in part enabled by the use of alpha-blending (see Eq.˜1): Since only the blended final color is used to compute the loss, individual primitives have no incentive to correctly model radiance.
Previous work [76] replaces the loss function of NeRF to apply the photometric loss to the individual color samples along the ray instead of the blended result, which effectively constrains such extreme cases. We find that naïvely applying this method to 3DGS significantly deteriorates reconstruction quality, because the vital D-SSIM loss term cannot be applied per-primitive. Instead, we introduce an additional weighted per-primitive $\mathcal{L}_{2}$-loss with the ground truth pixel color $\mathbf{I}$ to complement the standard 3DGS training loss [22]:
@@
${\mathcal{L}}_{\text{normal-var}}=\sum_{i=0}^{N-1}w_{i}(\bm{r})\left\lVert\bm{n}_{i}-\bm{N}\right\lVert_{2}^{2},$ (14)
@@ 0.A.2 Confidence
As $\hat{C}\rightarrow 0$, this quantity explodes; with our minimum confidence value, the gradient reaches at most 75-75 with $\beta=7.5\times 10^{-2}$. The maximum value was chosen to ensure that the gradients flowing back to individual Gaussians remain reasonably scaled, preventing further densification of already well-reconstructed regions. In addition, note that we detach $w_{i}(\bm{r})$ in Eq.˜11, as we do not want the Gaussians to modify their opacity guided by confidence gradients.
@@ 0.C.1 Color Variance Loss
$\frac{\partial{\mathcal{L}}_{\text{color-var}}}{\partial\text{sh}(\bm{\theta}_{i},\bm{d})}=2w_{i}(\bm{r})(\text{sh}(\bm{\theta}_{i},\bm{d})-\mathbf{I}).$ (22)
We can compute the gradient with regard to $\alpha_{i}$ (omitting $\bm{r}$ for notational convenience) as follows:
$\frac{\partial{\mathcal{L}}_{\text{color-var}}}{\partial w_{i}}$ $=\left\lVert\text{sh}(\bm{\theta}_{i},\bm{d})-\mathbf{I}\right\lVert_{2}^{2},$ (23)
$\frac{\partial w_{i}}{\partial\alpha_{i}}$ $=T_{i},$ (24)
$\frac{\partial{\mathcal{L}}_{\text{color-var}}}{\partial\alpha_{i}}=\frac{\partial{\mathcal{L}}_{\text{color-var}}}{\partial w_{i}}\frac{\partial w_{i}}{\partial\alpha_{i}}$ $=\left\lVert\text{sh}(\bm{\theta}_{i},\bm{d})-\mathbf{I}\right\lVert_{2}^{2}T_{i}.$ (25)
This gradient does not account for the effect of $\alpha_{i}$ on the Gaussians $j>i$, which include $(1-\alpha_{i})$ in $T_{j}$. Because $\alpha_{i}<1$ by definition, $\frac{\partial w_{j}}{\partial\alpha_{i}}=-\frac{w_{j}}{1-\alpha_{i}}$ holds and
$\sum_{j=i+1}^{N-1}\left\lVert\text{sh}(\bm{\theta}_{j},\bm{d})-\mathbf{I}\right\lVert_{2}^{2}\left(-\frac{w_{j}}{1-\alpha_{i}}\right)=-\frac{1}{1-\alpha_{i}}\sum_{j=i+1}^{N-1}w_{j}\left\lVert\text{sh}(\bm{\theta}_{j},\bm{d})-\mathbf{I}\right\lVert_{2}^{2}.$ (26)
@@
$\frac{\partial{\mathcal{L}}_{\text{color-var}}}{\partial\alpha_{i}}=\left\lVert\text{sh}(\bm{\theta}_{i},\bm{d})-\mathbf{I}\right\lVert_{2}^{2}T_{i}-\frac{1}{1-\alpha_{i}}\sum_{j=i+1}^{N-1}w_{j}\left\lVert\text{sh}(\bm{\theta}_{j},\bm{d})-\mathbf{I}\right\lVert_{2}^{2}$ (27)
For efficiency, $\sum_{j=i+1}^{N-1}w_{j}(\bm{r})\left\lVert\text{sh}(\bm{\theta}_{j},\bm{d})-\mathbf{I}\right\lVert_{2}^{2}$ is computed on-the-fly during the backward pass, via repeated subtraction from the final variance loss. Note that as we build upon SOF [52], we inherit the front-to-back backward pass from StopThePop [51].
@@ 0.C.2 Normal Variance Loss
$=\underbrace{\sum_{i=0}^{N-1}w_{i}(\bm{r})}_{=1-T}-2\underbrace{\left(\sum_{i=0}^{N-1}w_{i}(\bm{r})\bm{n}_{i}\right)\cdot\bm{N}}_{=\left\lVert\bm{N}\right\lVert_{2}^{2}}+\left\lVert\bm{N}\right\lVert_{2}^{2}\underbrace{\sum_{i=0}^{N-1}w_{i}(\bm{r})}_{=1-T}$ (30)
$=(1-T_{N})-(1+T_{N})\left\lVert\bm{N}\right\lVert_{2}^{2}.$ (31)
@@
$\frac{\partial{\mathcal{L}}_{\text{normal-var}}}{\partial\alpha_{i}}=\left\lVert\bm{n}_{i}-\bm{N}\right\lVert_{2}^{2}T_{i}-\frac{1}{1-\alpha_{i}}\sum_{j=i+1}^{N-1}w_{j}\left\lVert\bm{n}_{i}-\bm{N}\right\lVert_{2}^{2}$ (33)
@@ Appendix 0.E Appearance Embeddings
Appendix 0.F Mesh-Based Novel View Synthesis
Following MILo [18], we additionally provide numbers for mesh-based novel view synthesis, presented in Tab.˜9. We utilize the same script as open-sourced by MILo  [^7] https://github.com/Anttwo/MILo/blob/mesh_eval/milo/eval/mesh_nvs/render_mesh_nvs.py, ensuring a fair and reproducible comparison. For this purpose, we again utilize the Tanks & Temples dataset, as well as the Mip-NeRF 360 dataset [1]; we report results for all methods which extract unbounded meshes. Note that differently from NVS-focused methods such as 3DGS [28], we use all 6 scenes from the Tanks & Temples dataset.
As we can see, our method outperforms all other method for the Tanks & Temples dataset, whilst achieving the second-best scores across Mip-NeRF 360. Particularly for the Tanks & Temples dataset, our confidence-based losses and our improved appearance module directly translate to better geometry. For the Mip-NeRF 360 dataset, our method achieves the second-best scores for all metrics, although outperformed by MILo [18]. Note that as the coloring is achieved via a TensoRF backbone [4] (see MILo for details) conditioned on the rasterized depth, this evaluation setting does not favor methods with a higher primitive count; which would negatively impact MILo. Conversely, due to the inherent smoothness of the color field, this comparison benefits MILo, as it exhibits smoother geometry overall.
Table 9: Mesh-Based Novel View Synthesis: We report novel view synthesis metrics as a proxy for evaluating underlying geometric accuracy, using the Mip-NeRF 360 and Tanks & Temples datasets. We attain comparable quality to MILo [18] across both tested datasets.
Tanks & Temples Mip-NeRF 360
[…]
Ours 21.186 0.683 0.345 13.06M 23.639 0.661 0.332 18.43M
MILo 21.169 0.673 0.354 5.34M 23.869 0.675 0.322 6.20M
SOF 20.936 0.672 0.354 14.07M 23.304 0.634 0.351 23.28M
GOF 20.959 0.675 0.354 14.40M 22.998 0.619 0.365 22.37M